OpenStudy (anonymous):

Please help: Solve for x: 4/3 x + 2/3 = 6

OpenStudy (anonymous):

4/3x +2/3 = 6 4/3x = 6 - (2/3)n 4/3x = (16/3) x = (16/3)/(4/3) x = (16/3)*(3/4) x = 16/4 x = 4 i hope its clear

OpenStudy (anonymous):

4/3 x + 2/3 = 6\[\frac{ 4 }{ 3 }x + \frac{ 2 }{ 3 } = 6\]Subtract 2/3 from both sides. \[\frac{ 4 }{ 3 }x = 6 - \frac{ 2 }{3 }\]Multiply both sides by 3. \[4x = 18 - 2\]\[4x = 16\]Divide both sides by 4. \[x = 16/4\] What is x?

OpenStudy (anonymous):

4, thanks guys :)